Costs associated with the manufacture of miniature high-sensitivity piezoresistive pressure transducers is $76,000 per year. A clever industrial engineer found that by spending $16,000 now to reconfigure the production line and reprogram two of the robotic arms, the cost will go down to $58,000 next year and $50,000 in years 2 through 5-Using an interest rate of 8% per year, determine the present worth of the savings due to the reconfiguration. The present worth of the savings is determined to be $
Answer #1

Current Cost = $ 76000, Cost in Year 1 post innovation = $ 58000 and Cost in Year 2 post innovation = $ 50000

Year 1 Savings = 76000 - 58000 = $ 18000 and Year 2 to Year 5 Savings = 76000 - 50000 = $ 26000

Interest Rate = 8 % and Initial Investment = $ 16000

Present Worth = 18000 / (1.08) + 26000 / (1.08)^(2) + 26000 / (1.08)^(3) + 26000 / (1.08)^(4) 26000 / (1.08)^(5)= $ 96403.05